# Efficient evaluation of Bernstein-Bézier coefficients of B-spline basis functions over one knot span

### v1

@@ Abstract
New differential-recurrence relations for B-spline basis functions are given. Using these relations, a recursive method for finding the Bernstein-Bézier coefficients of B-spline basis functions over a single knot span is proposed. The algorithm works for any knot sequence which guarantees that all B-spline functions are at least $C^{0}$-continuous. It has good numerical behavior and has an asymptotically optimal computational complexity.
@@ 1 Introduction
$N_{0i}(u)=\left\{\begin{array}[]{ll}1&(u\in[t_{i},t_{i+1})),\\[4.30554pt] 0&\mbox{otherwise},\end{array}\right.$ (1.4)
@@
Let the adjusted Bernstein-Bézier basis form of the B-spline function $N_{mi}$ $(-m\leq i<n)$ over a single non-empty knot span $[t_{j},t_{j+1})\subset[t_{0},t_{n}]$ ($j=i,i+1,\ldots,i+m$) be
$N_{mi}(u)=\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad(t_{j}\leq u<t_{j+1}).$ (1.6)
@@
Find the adjusted Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$ $(0\leq k\leq m)$ (cf. (1.6)) of all non-trivial functions $N_{mi}$ over all non-empty knot spans $[t_{j},t_{j+1})\subset[t_{0},t_{n}]$, i.e., for $j=0,1,\ldots,n-1$ and $i=j-m,j-m+1,\ldots,j$ (cf. (1.5)).
@@
$N_{mi}(u)+\dfrac{t_{i}-u}{m}N_{mi}^{\prime}(u)=\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}\left(N_{m,i+1}(u)+\dfrac{t_{m+i+2}-u}{m}N_{m,i+1}^{\prime}(u)\right),$ (1.9)
@@
$N_{pq}\equiv 0\quad\mbox{for $q<-p$ or $q\geq n$}.$ (1.10)
@@ 2 The problem
At the end of our paper [7], we wrote that ”Further research is required to find the application of the new differential-recurrence relation in finding the Bernstein-Bézier coefficients of B-spline basis functions over a single knot span”. The point is that the proposed approach is too expensive if we want to compute the Bernstein-Bézier form of B-spline functions over one fixed knot span only. In short, the main reason is as follows: to find the Bernstein-Bézier coefficients of B-spline functions over non-empty interval $[t_{j},t_{j+1})$ using the recursive algorithm proposed by us, one has to previously compute the Bernstein-Bézier form of B-spline functions over all non-empty knot spans $[t_{n-1},t_{n}),[t_{n-2},t_{n-1}),\ldots,[t_{j+1},t_{j+2})$.
@@
Suppose the knot span $[t_{j},t_{j+1})$ is not empty with fixed $j$ $(0\leq j<n)$. Find the adjusted Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$ $(0\leq k\leq m)$ (cf. (1.6)) of all non-trivial functions $N_{mi}$ over $[t_{j},t_{j+1})$, i.e., for $i=j-m,j-m+1,\ldots,j$ (cf. (1.5)).
@@ 3 Some new results for B-spline basis functions
For any knot sequence (1.1), the B-spline functions $N_{mi}$ satisfy the following identities:
$mN_{mi}(u)+(t_{m+i+1}-u)N^{\prime}_{mi}(u)=m(t_{m+i+1}-t_{i})\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}},$ (3.1)
$mN_{mi}(u)+(t_{i}-u)N^{\prime}_{mi}(u)=m(t_{m+i+1}-t_{i})\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}},$ (3.2)
@@
Observe that Eqs (3.1) and (3.2) can be seen as fundamental properties of B-spline basis functions. Indeed, by adding (3.1) multiplied by $u-t_{i}$ and (3.2) multiplied by $t_{m+i+1}-u$ and next dividing by $m(t_{m+i+1}-t_{i})$, one can recover the famous the de Boor-Cox recurrence relation (see, e.g., [13, Eqs. (2.5)]), i.e.,
$N_{mi}(u)=(u-t_{i})\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}}+(t_{m+i+1}-u)\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}}\qquad(-m\leq i<n).$ (3.4)
We can also easily obtain the following well-known representation of the derivative of $N_{mi}$ in terms of two B-spline basis functions od degree $m-1$ (see, e.g., [13, Eqs. (2.7)]):
$N^{\prime}_{mi}(u)=m\left(\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}}-\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}}\right)\qquad(-m\leq i<n).$
@@
$N_{mi}(u)+\dfrac{t_{i}-u}{m}N_{mi}^{\prime}(u)=\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}\left(N_{m,i+1}(u)+\dfrac{t_{m+i+2}-u}{m}N_{m,i+1}^{\prime}(u)\right),$
@@
$N_{mi}(u)=\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad(t_{j}\leq u<t_{j+1}),$
@@ 4 The solution
If $m=0$ the problem is trivial, because $b^{(i,j)}_{0,0}=\delta_{ij}$ (cf. (1.4)). Here $\delta_{ij}$ denotes the Kronecker delta, i.e.,
@@ 4.1 The main idea
It means that to solve Problem 2.1 using Theorem 3.5 in $O(m^{2})$ time — i.e., to propose the asymptotically optimal algorithm of computing the adjusted Bernstein-Bézier forms of all non-trivial $N_{mi}$ over non-empty knot span $[t_{j},t_{j+1})$ for all $j-m\leq i\leq j$ — first, one needs to:
@@
    find the adjusted Bernstein-Bézier forms of $N_{m,j-m}$ and $N_{mj}$,
@@ 4.2 The initial conditions
$N_{mj}(u)=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+k}-t_{j})}B^{m}_{m}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad\qquad(t_{j}\leq u<t_{j+1}),$
@@
It means that the Bernstein-Bézier coefficients of $N_{m,j-m}$ and $N_{mj}$ $(m>0)$ in the non-empty knot span $[t_{j},t_{j+1})$ $(0\leq j\leq n-1)$ are as follows:
$b_{m0}^{(j-m,j)}=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+1}-t_{j-k+1})},\qquad b_{m,k}^{(j-m,j)}=0\quad(k=1,2,\ldots,m),$ (4.2)
@@
$b_{p,p}^{(i,j)}:=0\quad(i<j-p),\qquad b_{0,0}^{(i,j)}:=\delta_{ij},$ (4.9)
@@ 4.3 The algorithm
The $(m+1)^{2}$ adjusted Bernstein-Bézier coefficients $b_{m,k}^{(i,j)}$ (cf. (1.6)) of the basis B-spline functions $N_{mi}$ (1.2) over the knot span $[t_{j},t_{j+1})$, for $i=j-m,j-m+1,\ldots,j$ and $k=0,1,\ldots,m$, can be computed in the computational complexity $O(m^{2})$ in the following way:
@@
To examine the numerical behavior of the new algorithm (complexity $O(m^{2})$), we have compared it to the method mentioned in Remark 4.2 (complexity $O(m^{3})$). Numerical experiments show that both methods are stable and give almost the same results. The source code in C which was used to perform the tests is available at https://www.ii.uni.wroc.pl/~pwo/programs/BSpline-BF-one-interval.c.

### v2

@@ Abstract
New differential-recurrence relations for B-spline basis functions are given. Using these relations, a recursive method for finding the Bernstein-Bézier coefficients of B-spline basis functions over a single knot span is proposed. The algorithm works for any knot sequence and has an asymptotically optimal computational complexity. Numerical experiments show that the new method gives results which preserve a high number of digits when compared to an approach which uses the well-known de Boor-Cox formula.
@@ 1 Introduction
$N_{0,i}(u)=\left\{\begin{array}[]{ll}1&(u\in[t_{i},t_{i+1})),\\[4.30554pt] 0&\mbox{otherwise},\end{array}\right.$ (1.4)
@@
Let the adjusted Bernstein-Bézier basis form of the B-spline function $N_{m,i}$ $(-m\leq i<n)$ over a single non-empty knot span $[t_{j},t_{j+1})\subset[t_{0},t_{n}]$ ($j=i,i+1,\ldots,i+m$) be
$N_{m,i}(u)=\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad(t_{j}\leq u<t_{j+1}).$ (1.6)
@@
Find the adjusted Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$ $(0\leq k\leq m)$ (cf. (1.6)) of all non-trivial functions $N_{m,i}$ over all non-empty knot spans $[t_{j},t_{j+1})\subset[t_{0},t_{n}]$, i.e., for $j=0,1,\ldots,n-1$ and $i=j-m,j-m+1,\ldots,j$ (cf. (1.5)).
@@
$N_{m,i}(u)+\dfrac{t_{i}-u}{m}N_{m,i}^{\prime}(u)=\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}\left(N_{m,i+1}(u)+\dfrac{t_{m+i+2}-u}{m}N_{m,i+1}^{\prime}(u)\right),$ (1.9)
@@
$N_{p,q}\equiv 0\quad\mbox{for $q<-p$ or $q\geq n$}.$ (1.10)
@@ 2 The problem
At the end of our paper [7], we wrote that “Further research is required to find the application of the new differential-recurrence relation in finding the Bernstein-Bézier coefficients of B-spline basis functions over a single knot span”. The point is that the proposed approach is too expensive if we want to compute the Bernstein-Bézier form of B-spline functions over one fixed knot span only. In short, the main reason is as follows: to find the Bernstein-Bézier coefficients of B-spline functions over non-empty interval $[t_{j},t_{j+1})$ using the recursive algorithm proposed by us, one has to previously compute the Bernstein-Bézier form of B-spline functions over all non-empty knot spans $[t_{n-1},t_{n}),[t_{n-2},t_{n-1}),\ldots,[t_{j+1},t_{j+2})$.
@@
Suppose the knot span $[t_{j},t_{j+1})$ is not empty with fixed $j$ $(0\leq j<n)$. Find the adjusted Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$ $(0\leq k\leq m)$ (cf. (1.6)) of all non-trivial functions $N_{m,i}$ over $[t_{j},t_{j+1})$, i.e., for $i=j-m,j-m+1,\ldots,j$ (cf. (1.5)).
@@ 3 Some new results for B-spline basis functions
For any knot sequence (1.1), the B-spline functions $N_{m,i}$ satisfy the following identities:
$mN_{m,i}(u)+(t_{m+i+1}-u)N^{\prime}_{m,i}(u)=m(t_{m+i+1}-t_{i})\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}},$ (3.1)
$mN_{m,i}(u)+(t_{i}-u)N^{\prime}_{m,i}(u)=m(t_{m+i+1}-t_{i})\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}},$ (3.2)
@@
Observe that Eqs (3.1) and (3.2) can be seen as fundamental properties of B-spline basis functions. Indeed, by adding (3.1) multiplied by $u-t_{i}$ and (3.2) multiplied by $t_{m+i+1}-u$ and next dividing by $m(t_{m+i+1}-t_{i})$, one can recover the famous de Boor-Cox recurrence relation (see, e.g., [13, Eqs. (2.5)]), i.e.,
$N_{m,i}(u)=(u-t_{i})\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}}+(t_{m+i+1}-u)\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}}\qquad(-m\leq i<n).$ (3.4)
We can also easily obtain the following well-known representation of the derivative of $N_{m,i}$ in terms of two B-spline basis functions of degree $m-1$ (see, e.g., [13, Eqs. (2.7)]):
$N^{\prime}_{m,i}(u)=m\left(\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}}-\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}}\right)\qquad(-m\leq i<n).$
@@
$N_{m,i}(u)+\dfrac{t_{i}-u}{m}N_{m,i}^{\prime}(u)=\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}\left(N_{m,i+1}(u)+\dfrac{t_{m+i+2}-u}{m}N_{m,i+1}^{\prime}(u)\right),$
@@
If an inner knot $t_{i}$ ($i=1,2,\ldots,n-1$) has multiplicity greater than $m$, then every B-spline function of degree $m$ based on these knots is identical to zero either for $t>t_{i}$ or for $t<t_{i}$. Then the B-spline basis consists of two independent bases and the problem can be divided into two smaller ones: one for the non-empty knot spans in the interval $(t_{0},t_{i})$ and the other in $(t_{i},t_{n})$.
@@
$N_{m,i}(u)=\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad(t_{j}\leq u<t_{j+1}),$
@@ 4 The solution
If $m=0$ the problem is trivial, because $b^{(i,j)}_{0,0}=\delta_{i,j}$ (cf. (1.4)). Here $\delta_{i,j}$ denotes the Kronecker delta, i.e.,
@@ 4.1 The main idea
It means that to solve Problem 2.1 using Theorem 3.5 in $O(m^{2})$ time — i.e., to propose the asymptotically optimal algorithm of computing the adjusted Bernstein-Bézier forms of all non-trivial $N_{m,i}$ over non-empty knot span $[t_{j},t_{j+1})$ for all $j-m\leq i\leq j$ — first, one needs to:
@@
    find the adjusted Bernstein-Bézier forms of $N_{m,j-m}$ and $N_{m,j}$,
@@ 4.2 The initial conditions
$N_{m,j}(u)=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+k}-t_{j})}B^{m}_{m}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad\qquad(t_{j}\leq u<t_{j+1}),$
@@
It means that the Bernstein-Bézier coefficients of $N_{m,j-m}$ and $N_{m,j}$ $(m>0)$ in the non-empty knot span $[t_{j},t_{j+1})$ $(0\leq j\leq n-1)$ are as follows:
$b_{m,0}^{(j-m,j)}=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+1}-t_{j-k+1})},\qquad b_{m,k}^{(j-m,j)}=0\quad(k=1,2,\ldots,m),$ (4.2)
@@
$b_{p,p}^{(i,j)}:=0\quad(i<j-p),\qquad b_{0,0}^{(i,j)}:=\delta_{i,j},$ (4.9)
@@ 4.3 The algorithm
The $(m+1)^{2}$ adjusted Bernstein-Bézier coefficients $b_{m,k}^{(i,j)}$ (cf. (1.6)) of the basis B-spline functions $N_{m,i}$ (1.2) over the knot span $[t_{j},t_{j+1})$, for $i=j-m,j-m+1,\ldots,j$ and $k=0,1,\ldots,m$, can be computed in the computational complexity $O(m^{2})$ in the following way:
@@
To examine the numerical behavior of the new algorithm (complexity $O(m^{2})$), we have compared it to the method mentioned in Remark 4.2 (complexity $O(m^{3})$). Numerical experiments show that both methods are stable and give almost the same results. As expected, the new algorithm is usually faster, especially for larger $m$. The source code in C which was used to perform the tests is available at https://github.com/filipchudy/bspline-coeffs-one-interval. The results have been obtained on a computer with Intel Core i5-6300U CPU at 2.40GHz processor and 8GB RAM, using GNU C Compiler 13.2.0 (double precision).
Example 4.4.
An experiment which checks the numerical quality of the method given in [7] and the new algorithm has been done. For $m\in\{3,4,5,10,20,30,50\}$ and $n\in\{10,50,100\}$, the following procedure has been repeated 50,000 times.
A sequence of knots has been generated in the following way. A sequence of knots has been randomized using the rand C function. The first knot is sampled from the $[-10,10]$ interval. Then, the following knots are sampled, with randomized multiplicities and distances, from $1,2,\ldots,m$ and $(0,0.5)$, respectively. Afterwards, all right-hand side boundary knots are clamped.
Then, all Bernstein-Bézier coefficients of the B-spline functions are evaluated using the method given in [7], the new algorithm, and the approach given in Remark 4.2. The evaluation has been done for each knot span. Table 4.1 shows the mean number of correct decimal digits for the first two methods, with reference results being the ones obtained by the third method, which uses the de Boor-Cox formula. The mean number is over all $(m+1)^{2}$ Bernstein-Bézier coefficients of B-spline functions in each knot span, over all non-empty knot spans, for 50,000 different knot sequences.
$m$ $n$ mean [7] mean new
3 10 13.930 17.858
3 50 13.952 17.859
3 100 13.959 17.860
4 10 13.895 17.762
4 50 13.902 17.763
4 100 13.906 17.764
5 10 13.817 17.650
5 50 13.805 17.653
5 100 13.804 17.654
10 10 13.533 17.310
10 50 13.402 17.317
10 100 13.382 17.319
20 10 13.642 17.031
20 50 13.072 16.943
20 100 12.985 16.946
30 10 13.792 16.905
30 50 12.861 16.646
30 100 12.706 16.651
50 10 13.903 16.775
50 50 12.532 16.137
50 100 12.303 16.150
Table 4.1: The mean number of correct digits for Example 4.4. The source code in C which was used to perform the tests is available at https://github.com/filipchudy/bspline-coeffs-one-interval.
Example 4.5.
A comparison of running times for the new algorithm and the method given in Remark 4.2 has been done. For $m\in\{3,4,5,10,20,30,50\}$ and $n\in\{10,50,100\}$, the following procedure has been repeated 50,000 times.
A sequence of knots has been generated in the following way. A sequence of knots has been randomized using the rand C function. The first knot is sampled from the $[-10,10]$ interval. Then, the following knots are sampled, with randomized multiplicities and distances, from $1,2,\ldots,m$ and $(0,0.5)$, respectively. Unlike Example 4.4, the right-hand side boundary knots are not clamped.
Then, all Bernstein-Bézier coefficients of the B-spline functions over all non-empty knot spans are evaluated using the new algorithm and the method given in Remark 4.2. Table 4.2 shows the comparison of running times in seconds, for different $m,n$, for all 50,000 tests.
$m$ $n$ new deBoor
3 10 0.158 0.180
3 50 0.523 0.643
3 100 0.962 1.206
4 10 0.190 0.238
4 50 0.668 0.921
4 100 1.261 1.770
5 10 0.226 0.309
5 50 0.843 1.270
5 100 1.602 2.448
10 10 0.404 0.847
10 50 1.730 3.916
10 100 3.377 7.742
20 10 0.746 2.670
20 50 3.561 13.452
20 100 7.152 27.334
30 10 1.099 5.635
30 50 5.397 28.847
30 100 10.766 57.869
50 10 1.811 14.814
50 50 9.176 77.182
50 100 18.226 153.975
Table 4.2: Running times comparison (in seconds) for Example 4.5. The source code in C which was used to perform the tests is available at https://github.com/filipchudy/bspline-coeffs-one-interval.